KR’96

Fifth International Conference on Principles of Knowledge Representation and Reasoning

KR’96

Cambridge, Massachusetts, U.S.A.  World Wide Web: http://www.kr.org/kr/
November 5–8, 1996  Autoresponder: kr96-info@kr.org

Explicit representations of knowledge manipulated by inference algorithms provide an important foundation for much work in Artificial Intelligence, from natural language dialogue systems to expert systems. We intend KR’96 to be a place for the exchange of news, issues, and results among the community of researchers in the principles and practices of knowledge representation and reasoning (KR&R) systems.

We encourage papers that present substantial new results in the principles of KR&R systems while clearly showing the applicability of those results to implemented or implementable AI systems. We also encourage “reports from the field” of applications, experiments, developments, and tests. The following topics are meant to be suggestive of the scope of the conference.

Representational Formalisms Representations of Belief,

Intention Time, Space, Action, Events

Nonmonotonic Logics Description Logics

Reasoning Techniques Deduction Induction Abduction Reasoning under Uncertainty Parallel and Distributed Implementations Efficiency Measures and Complexity Implemented KR&R Systems

Reports, Updates

Comparisons, Evaluations

Significant Applications Planning Robotics Diagnosis Natural Language Multi-Agent Environments Knowledge Bases

Implications for/of Machine Learning Decision Theory Databases Software Engineering

SCHEDULE

May 6, 1996  Extended abstracts due
July 1, 1996  Results to authors
August 15, 1996  Final papers due
November 5-8, 1996 KR’96

KR’96 will be held in Cambridge immediately preceding the AAAI Fall Symposia Series, and immediately after several independent workshops. Tentative information for the adjoining meetings is as follows, with all located in Cambridge, Massachusetts.

The AAAI Fall Symposia Series will be held November 9–11, 1996. For more information,

see http://www.aaai.org/.

Description Logic ’96 will be held November 2–4, 1996. For more information contact the organizing committee at dl96@dl.kr.org. The organizing committee consists of Lin Padgham (chair), Deborah McGuinness, Peter Patel-Schneider, Enrico Franconi, and Manfred Gehrke.

Relevance in Knowledge Representation and Reasoning (RRR-96) will be held November 2–4, 1996. For more information, contact the organizers, Alon Levy and Russ Greiner, at

levy@research.att.com and greiner@scr.siemens.com.

SUBMISSION OF PAPERS: The Program Committee will review EXTENDED ABSTRACTS rather than complete papers. Submissions must be at most twelve (12) pages, excluding the title page and the bibliography, with a maximum of 38 lines per page and an average of 75 characters per line (corresponding to the LaTeX article-style, 12pt). Overlength submissions will be rejected without review. All abstracts must be submitted on 8 1/2 by 11 inch or A4 paper, and printed or typed in 12-point font (10 characters per inch on a typewriter). Dot matrix printout, FAX, or electronic submission will not be accepted. Each submission should include the names and complete addresses (including email, when possible) of all authors. Correspondence will be sent to the first author, unless otherwise indicated. Also, authors should indicate under the title which of the topic areas listed above best describes their paper (if none is appropriate, please give a set of keywords that best describe the topic of the paper).

KR’96 is arranging with AAAI to handle the collection and acknowledgment of submissions. To be considered, five (5) paper copies of each extended abstract must be received no later than May 6, 1996 at the following address:

KR’96 c/o AAAI 445 Burgess Drive Menlo Park, CA 94025

Receipt of submissions will be acknowledged, ordinarily by email. Remaining questions concerning receipt of submission may be addressed to AAAI at

Tel: 415-328-3123 Fax: 415-321-4457 Email: kr@aaai.org

MULTIPLE SUBMISSIONS: Submitted papers must be unpublished and substantively different from papers currently under review.

NOTIFICATION: Authors will be notified of the Program Committee’s decision by July 1, 1996. Notification will be made by electronic mail whenever possible.

FINAL PAPERS: Authors of accepted papers will be expected to submit substantially longer full papers for the conference proceedings. Final camera-ready copies of the full papers will be due August 15, 1996. Final papers will be allowed at most twelve (12) double-column pages in the conference proceedings (corresponding to approximately 28 article-style LaTeX pages; a style file will be provided by the publisher).

REGISTRATION: Registration, lodging, and travel information will be distributed later; check the web page or autoresponder listed above for current information. KR’96 is arranging with AAAI to handle registration, including payment by credit card.

CONFERENCE CHAIR Jon Doyle MIT Laboratory for Computer Science 545 Technology Square Cambridge, MA 02139 USA Voice: +1 (617) 253-3512 Fax: +1 (617) 258-8682 EMAIL: doyle@mit.edu

PROGRAM CO-CHAIRS

Luigia Carlucci Aiello

Universit di Roma La Sapienza

Dipartimento di Informatica e Sistemistica via Salaria 113 00198 Roma ITALY Voice: +39 6 8841947 Fax: +39 6 85300849 EMAIL: aiello@dis.uniroma1.it Stuart C. Shapiro State University of New York at Buffalo Department of Computer Science 226 Bell Hall Buffalo, NY 14260-2000 USA Voice: +1 716 645 3180 ext. 125 Fax: +1 716 645 3464 EMAIL: shapiro@cs.buffalo.edu

INTER-CONFERENCE COOPERATION CHAIR

Ronald P. Loui

Washington University, USA EMAIL: loui@cs.wustl.edu

PUBLICITY CHAIR

Werner Horn

Austrian Research Institute for Artificial Intelligence Austria EMAIL: werner@ai.univie.ac.at

PROGRAM COMMITTEE (Preliminary)

Syed Ali (SW. MO St. U., USA) Fahiem Bacchus (U. Waterloo, Canada) Afzal Ballim (EPFL, CH) John A. Barnden (NM St. U., USA) Ron Brachman (AT&T Bell Labs, USA) Maurice Bruynooghe (Catholic Univ. of Leuven, BE) Anthony G. Cohn (U. Leeds, UK) Marie Odile Cordier (IRISA, FR) Ernest Davis (NYU, USA) Didier Dubois (IRIT, FR) Thomas Eiter (T.U. Wien, AT) Luis Farinas del Cerro (IRIT, FR) Richard Fikes (Stanford U., USA) Dov Gabbay (Imperial College, UK) Peter Gaerdenfors (Lund U., SE) Mike Georgeff (AAII, AU) Fausto Giunchiglia (U. Trento, Italy) Frank van Harmelen (Free Univ. of Amsterdam, NL) Patrick Hayes (U. IL, USA) Jim Hendler (U. Md, USA) Eduard Hovy (USC/ISI, USA) Hiroachi Kitano (Sony, JP) Kurt Konolige (SRI, USA) Sarit Kraus (Bar Ilan U., IL) David Israel (SRI, USA)

Lucja Iwanska (Wayne St. U., USA) Benjamin Kuipers (U. TX, USA) Deepak Kumar (Bryn Mawr Coll., USA) Gerhard Lakemeyer (U. Bonn, Germany) Fritz Lehmann (Cycorp and GRANDAI, USA) Doug Lenat (Cycorp, USA) Maurizio Lenzerini (U. Roma, IT) Hector Levesque (U. Toronto, Canada) Vladimir Lifschitz (U. TX, USA) Robert MacGregor (USC/ISI, USA) Joao Martins (Tech Univ of Lisbon, PT) Riichiro Mizoguchi (U. Osaka, JP) Bernhard Nebel (U. Ulm, DE) Hwee Tou Ng (DSO, Singapore) Hans Juergen Ohlbach (Imperial College, UK) Lin Padgham (RMIT, AU) Ramesh Patil (USC/ISI, USA) Anand Rao (AAII, Australia) Ray Reiter (U. Toronto, Canada) Jeff Rosenschein (Hebrew U., IL) Erik Sandewall (Linkoeping U., SE) Len Schubert (U. Rochester, USA) John Sowa (U. Binghamton, USA) Piero Torasso (U. Torino, IT) Wolfgang Wahlster (DFKI, Germany)